

********************************************************************
UARS WEEKLY STATUS REPORT - 11 FEB 2000
********************************************************************
***********
SOLSTICE
***********
After ending daylight only operations on Tuesday 2/8, we began to again
gather solar data on all available orbital tracking periods and stellar
calibration observations whenever possible. The reduced data return has not
prevented us from meeting all the SOSLTICE science objectives.
***********
SUSIM
***********
SUSIM continues to gather a minimum set of daily solar scans scheduled
for periods of predicted telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM spectrometers continue to be active under PMON control*. While
under PMON control, the PEM spectrometers computed energy deposition from
electrons and protons measured in-situ by the particle spectrometers.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
MLS continues to operate in a mode where limb scans are performed on the day
side of the orbit at ~50N-80N latitudes, with the 205 GHz radiometer on for
measurements of ClO, HNO3 and O3.
All electronics that are on continue to work flawlessly, and no scan slips have
occurred this observing period (since the 31 January 2000 turn on).
Unfortunately, a large fraction of the good limb scans are being lost due to
gaps in the UARS telemetry stream.
MLS is now scheduled to remain on until ~13 February.
The arctic data are being processed at JPL, and the data mapped to the 465 K
potential temperature surface (lower stratosphere) are available from the MLS
web site (http://mls.jpl.nasa.gov) by clicking on "UARS MLS Data" and then on
"UARS MLS arctic observations for 1-10 February 2000."
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MLS instrument team asked for additional days of MLS
operation. Data being received were of excellent quality and
measurements were being made in the 50-80 degree latitude range.
They requested additional days of MLS operation, extending as
long as spacecraft power is adequate for the instrument until the
yaw maneuver on 15 February. To provide adequate spacecraft
power for MLS, it was necessary to reschedule HRDI to day-only
observations, and to turn HALOE off. The MPG, along with the
UARS Project Scientist, contacted HRDI and HALOE instrument
personnel so that appropriate changes HRDI and HALOE Daily
Activity Plans could be made.
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
18 days WIN (5.11A) PROD (L1-L3) (520R-391C-14F=115L)
0 day SOL (4.62) PROD (L1) (9R-0C=9L)
3 days SEN (5.0) PROD (1992) (3R-3C=0L)
The CDHF Continus to process 1992 SENSOR data as requested by FDF.
CDHF REHOST:
- Successfully processed ATTITUDE EXTRACTION V7.3 Software on
Rehost System against LEVEL0 data (DAY 3063) ingested from the LZP.
- Successfully tested SUSIM V7.2 Patch delivery on Rehost
System. processed same designated day on old system to use as a
comparison. SUSIM has validated.
***********
FOT
***********
REPORTING PERIOD
This report covers 29 January, 2000 (Orbit 45814, GMT Day 029) through 4
February 2000 (Orbit 45918, GMT Day 035).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit
HRDI 029/0836 45734
MLS 031/1807 45734
ACRIM 034/0107 45743
SUSIM 034/0107 45753
HALOE 034/1409 45763
The beta angle ranged between the angles of 74.8 degrees to 52.1 degrees for
this report period. The beta angle is now decreasing toward a minimum angle of
3.0 degrees on 15 February 1999 (DOY 046).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.0 to 1.3 degrees C. Additional battery performance data for each
battery on 28 January 2000 and 04 Febraury 2000 are provided in Appendix D.
The clock error ranged from -13.2 to 6.1 msec during this report period. Clock
rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% OUT-OF-LIMITS = 04
% ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 05
% MLS = 02
% HRDI = 01
% ACRIM = 01
% SUSIM = 01
% HALOE = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 03
Two (2) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 00/021 - 2287.496849
Transponder B: 00/031 - 2287.498836
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/034 show the observatory data loss to be 565 hours, 15
minutes, 06.832 seconds (an increase of 5 minutes 6.832 seconds since the last
report period).
The increase of 5 minutes 6.832 seconds occurred over three (3) days:
01 minutes 17.424 seconds on Day 028 due to an unspecified problem.
03 minutes 08.448 seconds on Day 029 due to an unspecified problem.
00 minutes 40.960 seconds on Day 030 due to an unspecified problem.
These calculations are under repair. They are being recalculated from the date
where PASS 2 operations were scheduled exclusively.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
Nominal operations only this report period.
MEETINGS
2 February, 2000 (DOY 033) - The Monthly meeting was held. Charlie Jackman
contacted all the instrument PI's and confirmed that all wanted to go to 3
TDRSS operations, if the risk to UARS was less than 10%.
SUSIM was asked whether there quarterly calibrations could be broken into
smaller pieces. They didn't think it could be done but they are looking into
it. This leads to more discussion about using the recorder on an as needed
basis instead of continually like the current operations.
4 February, 2000 (DOY 035) - A meeting took place to introduce the new
simulator sustaining engineer, Adrian Suskind, to the FOT and the facility.
She is a part of code 582 and will support the simulator entirely.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for February 15, 2000 (DOY 046). This
maneuver will be a Reverse to Forward Flight maneuver (Southern Hemisphere
viewing).
A partial solar eclipse of sun will occur on Saturday, February 5, 2000 (DOY
036). The maximum shadow over two orbits will be approximately 40%. No impact
to operations is expected to occur.
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 4 February 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
45819 029/0836 00-026 HRDI Safehold - Hi TBAT
45849 031/0926 00-032 OBC Clk Adj 41.2 to 20.6 msec/day
45857 031/1807 00-033 MLS Power ON
45871 032/1949 00-034 MLS Microload
45880 033/1046 00-035 NBTR NBTR NOOPS
45889 034/0107 00-036 ACRIM Power OFF
45889 034/0107 00-037 SUSIM Power OFF
45895 034/0934 00-038 OBC Clk Adj 20.6 to 0.00 msec/day
45897 034/1409 00-039 HALOE Power ON
45901 034/1914 00-040 NBTR NBTR NOOPS
45915 035/1817 00-041 NBTR NBTR NOOPS
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 4 February 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPTSBARTMP Y-LO HRDI Barrel Temp 45815-16,20-21,865,66
HRPTSBARTMP R-LO HRDI Barrel Temp 45817-19
QTZEPSIF2 Y-LO ZEPS I/F Temp Fwd 45845-51
QTZEPSIF1 Y-LO ZEPS I/F Temp Aft 45845-51
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
No Alerts to report this report period.
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 4 February 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
45849 31/09:30 (31 Jan 00) from 41.2 to 20.6 msec/day
45895 034/09:45 (03 Feb 00) from 20.6 to 41.2 msec/day
_______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 4 February 2000
28 Jan 2000 - GMT Day 028 Beta = 74.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- -------- --------- ------- ------ ------- ------
1 -67.2/-89.6 -2.42/-3.68 +0.0/+0.0 16.2 0.000 0.0
2 +0.0/-33.6 +2.92/+1.66 +16.8/-14.8 27.5 1.057 14.0
3 +44.8/-5.6 +1.97/+1.03 +16.8/-14.4 27.5 1.061 14.5
04 Feb 2000 - GMT Day 035 Beta = 52.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ----- ------- ------
1 -67.2/-89.6 -3.05/-4.62 +0.0/+0.0 16.0 0.000 0.0
2 +11.2/-28.0 +5.47/+4.51 +26.8/-16.0 25.8 1.040 21.1
3 +67.2/-5.6 +4.51/+3.23 +26.0/-16.0 25.8 1.038 20.8
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 4 February 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-001 FOT 45816 029/0416 14170 Generic Late Acq
00-002 FOT 45820 029/1022 21919 TLM Dropout
00-001 FOT 45840 030/1902 14170 Generic Late Acq
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS-
- ----- -------- ----- ---- --- -----------------
00-001 FOT 45816 029/041 14170 Generic Late Acq
00-002 FOT 45820 029/10 21919 TLM Dropout
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
No new reports generated
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-064 FOT 44135 10/8-9/99 N/A HRDI Cal Lamp Errors
99-071 FOT 44398 299/1212 21512 TPF Not Receiving 512
PB
SEVNTFW00/05